# Training Strategies for Isolated Sign Language Recognition

### v1

@@ 4 Datasets
We assess the training strategies described above on three large-scale ISLR datasets: WLASL [25], AUTSL [37], and Slovo [23]. They are the most diverse open-source data in terms of dataset contributors, performing gestures (signers), and contexts (see Table 1). Thus, the evaluation only with them is sufficient to confirm the pipeline’s effectiveness.
@@
  • Slovo is the largest, most diversified, and the only publicly available Russian Sign Language (RSL) dataset. The dataset contains 20,400 HD and FHD untrimmed videos performed by 194 signers. It is divided into 1,001 classes, including the additional “no event" class, which indicates videos where the signer is not performing a sign. The Slovo was produced by crowdsourcing platforms, providing lifelike backgrounds and lighting conditions. RSL also contains non-invariant signs to horizontal flip.
@@ 4.1 SlovoExt Dataset
We expand the Slovo dataset with a self-assembled 30,600 videos. The combination is called SlovoExt and comprises 51,000 samples divided into the same 1,001 classes as Slovo (see Table 1). The process of SlovoExt creation is constructed similarly to maintain the Slovo distribution characteristics such as video length, signer’s appearance, and context heterogeneity identical (see Appendix B). SlovoExt is also utilized to evaluate the proposed training strategies.
The dataset was created exclusively by native speakers of RSL and interpreters proficient in RSL. We involved diverse contributors in the data collection process to address concerns about potential differences in sign presentation between deaf and hard of hearing people. This approach aims to help neural networks manage gesture variability and mitigate the effects of native sign bias [11].
@@
Table 2: Evaluation results. We present two top-1 accuracy metrics for each setup: the metric on the basic approach and the metric obtained via the proposed training pipeline. Two pre-trained models on the K400 dataset [24] are utilized: MViTv2 trained on the classification task and MaskFeat trained in a self-supervised manner to reconstruct masked pixels.
@@ 5 Experiments
We assess the effectiveness of the training pipeline by comparing two metrics: utilizing (1) the basic approach [^4] The basic approach is the same as the proposed pipeline, but without video and image augmentations, regression head, and classification loss balancing. and (2) proposed strategies. We fine-tune three architectures: two transformers (MViTv2-S with different pre-training and MViTv2-B [26]) and one CNN (I3D [5] with a ResNet-50 backbone) — on four ISLR datasets. The experiment results are provided in Table 2.
@@ 5.1 Preprocessing
Input videos are resized to 300 resolution and converted to HDF5 video format. During the training stage, videos are square-padded and randomly cropped to achieve a resolution of 224 × 224. The videos’ low side is resized to 224 and square-padded during the validation and testing stages. The model input is a sequence of 32 frames sampled with a step of 2. If the video does not have enough frames for sampling, the last frame is repeated to form a complete clip. The Slovo and the SlovoExt combination are subjected to excluding non-mirrored signs from horizontal flip transformation being applied to them. The WLASL signs are not being flipped, and the AUTSL signs are flipped with no exceptions because Section 6 shows the efficacy of such decisions.
@@
Table 3: Ablation study results. We employ the MViTv2-S model to assess the influence of each component by removing it from the pipeline and comparing the results. Accuracy is evaluated on the SlovoExt combination.
@@
5.2 Training Strategies Application
As shown in Figure 2, the preprocessed video is initially exposed to video augmentations. One randomly selected from four modifications (“speed up", “slow down", “random add", and “random drop") is applied to each batch. In contrast, “random boundary shift" is constantly affected. In experiments, “speed up" increases the video speed by 2 times, and “slow down" decreases it by 2 times. “Random drop" shortens video by 10%, while “random add" lengthen it by 30%. The random values for sign boundaries shifting are from the interval [-5, 5], except the interval [-5, 0] for the WLASL and AUTSL datasets due to their trimming by the sign boundaries.
Image augmentations modify the video batch by randomly choosing a combination of them with expertly selected probabilities and magnitudes (see Table 6 in Appendix). We employ Downscale and Image Compression realizations from the Albumentations [3] computer vision tool and Random Noise, Sharpness, Horizontal Flip, and Color Jitter from the MMAction2 [8] open-source toolbox for video understanding. Note that CutMix [41] and MixUp [43] influence only transformer models in our experiments because they can degrade CNNs due to a lack of correlation between neighboring pixels [2].
@@ 5.2 Training Strategies Application
5.3 Training Details
Training is performed until convergence on four Tesla H100s with 80GB RAM. Early stopping is triggered if the top-1 accuracy metric does not increase by at least 0.003 after 7 epochs. For the first 20 epochs, the learning rate is modified by a linear scheduler, and a cosine scheduler is used from epochs 20 to 100. Cross-entropy loss is minimized using the AdamW optimizer. Other parameters are variable and described in Table 7.
@@ 5.4 Results
As Table 2 shows, the proposed pipeline consistently improves top-1 accuracy across different datasets and architectures. The impact is less prominent but still noticeable regarding the model I3D, resulting in an average improvement of 1.72% across datasets.
The developed strategies allow the achievement of state-of-the-art results on the WLASL and Slovo datasets. These metrics are obtained using MViTv2-S pre-trained with MaskFeat. We utilize the additional pretraining on the SlovoExt dataset for WLASL, resulting in a 62.89% top-1 accuracy. The state-of-the-art result on the Slovo is also achieved with K400-MViTv2 initialization, surpassing the current one by 13.06% mean accuracy instead of 14.12% with MaskFeat.
@@ 6 Ablation Study
This section estimates the impact of each part of the proposed pipeline individually. We divide all modifications into three blocks: (1) image augmentations, (2) video augmentations, and (3) additional losses. We test the necessity for each block by disconnecting it from the entire pipeline and for its parts by shutting down only a particular change and fixing everything else. The ablation study uses MViTv2-S pre-trained on the MaskFeat to train on the SlovoExt combination.
@@ 6.1 Image Augmentations Necessity
The image augmentations block is additionally divided into two parts – basic image transformations and CutMix-MixUp pair – to simplify assessing their influence. We shut down each part and a combination of them, resulting in three experiments. The first three rows in Table 3 illustrate that the metric significantly decreases by 4.06% in the absence of all image augmentations. The influence of basic image augmentation is two times less than the CutMix-MixUp one.
We also evaluate various augmentation-applying approaches, including RandAugment [9], UniformAugment [28], and random sequential applications. Table 4 indicates that RandAugment and UniformAugment approaches are not justified in the proposed pipeline.
@@ 6.2 Video Augmentations Necessity
Similar to the above subsection, we split the estimation of video augmentations necessity by disconnecting each of them separately and the whole block, providing six experiments. The most notable impact achievable with one modification is produced by “slow down" augmentation, as without it, the top-1 accuracy drops by 0.66%.
@@ 7 Discussion
Positional Statement. Through the research of the ISLR task, we involved the All-Russian Community of Deaf experts and professional sign language interpreters. The expertise of the All-Russian Society of the Deaf [^7] https://voginfo.ru/all-russian-society-of-the-deaf/ was utilized at every stage of the SlovoExt dataset creation, including data collection, validation, and verification processes, as well as video tagging. We also involved deaf consultants in developing training strategies to apply Considerations to particular solutions. Additionally, some of our researchers took formal courses on RSL to enhance their knowledge of this domain.
Limitations. Since the failure case of CNNs metrics degradation in Table 2 because of CutMix and MixUp, such augmentations must be excluded from the proposed pipeline while learning CNNs. Besides, the regression head and IoU-balanced CE loss require sign boundary annotations that are frequently difficult to achieve. In another situation, the dataset may be trimmed by sign boundaries. Thus, “random boundary shift" augmentation can shift the sign boundaries only inward to the center instead of the entire application with a shift to both sides. We processed RGB frames and did not analyze articulation, keypoints, and depth information. The lack of awareness about non-mirrored signs in sign language can be problematic for researchers with flip augmentation. There are other hardships in each sign language, and each has its specifics that are important to consider when adapting the pipeline.
@@ 8 Conclusion
In this paper, we introduce a training pipeline for the ISLR models, considering the specifics of the SLR domain and the constraints of real-world usage. We prove the effectiveness of applying image and video augmentations to address the issues of low data quality and varying gesturing speed. The importance of integrating temporal information into the model by the regression head combined with IoU-balanced CE loss is also presented. The developed pipeline delivers state-of-the-art results on the WLASL and Slovo datasets. Future work will extend the training strategies to CSLR and SLT tasks.
@@ Appendix B SlovoExt Creation Pipeline
We collected the self-assembled part of the SlovoExt dataset in three main stages: (1) data mining, (2) validation, and (3) annotation. Tasks were available only to those who passed the mandatory Russian Sign Language (RSL) exam and could thus be considered native speakers. We utilized Yandex Toloka [^8] https://platform.toloka.ai/ for the first stage and ABC Elementary [^9] https://elementary.activebc.ru for the others due to the specifics of each of the platforms and to avoid checking signers own videos. The creation pipeline of the self-assembled Russian Isolated Sign Language Dataset is briefly described below.

### v2

@@ 4 Datasets
We assess the training strategies described above on three large-scale ISLR datasets: WLASL [23], AUTSL [36], and Slovo [21]. They are the most diverse open-source data in terms of dataset contributors, performing gestures (signers), and contexts (see Table 1). Evaluating on these datasets alone is sufficient to confirm the pipeline’s effectiveness.
@@
  • Slovo is the largest, most diversified, and the only publicly available RSL dataset. The dataset contains 20,400 HD and FHD untrimmed videos performed by 194 signers. It is divided into 1,001 classes, including the additional “no event" class, which indicates videos where the signer is not performing a sign. Slovo was crowdsourced, featuring real-life conditions and non-invariant signs to horizontal flip.
@@ 4.1 SlovoExt Dataset
We expand the Slovo dataset with a self-assembled 30,600 videos. The combination is called SlovoExt and comprises 51,000 samples divided into the same 1,001 classes as Slovo (see Table 1). The process of SlovoExt creation is constructed similarly to maintain the Slovo distribution characteristics such as video length, signer’s appearance, and context heterogeneity identical. SlovoExt is also utilized to evaluate the proposed training strategies.
The dataset was created by native speakers of RSL and interpreters proficient in RSL. We involved diverse contributors in the data collection process to address concerns about differences in sign presentation between deaf and hard of hearing people. This approach aims to help neural networks manage gesture variability and mitigate the effects of native signer bias [9].
@@
Table 2: Evaluation results. We present two top-1 accuracy metrics for each setup: the metric on the basic approach and the metric obtained via the proposed training pipeline. Two pre-trained models on the K400 dataset [22] are utilized: MViTv2 trained on the classification task and MaskFeat trained in a self-supervised manner to reconstruct masked pixels. Green values show gains over the basic approach.
@@ 5 Experiments
The effectiveness of the training pipeline is assessed by comparing two metrics: utilizing (1) the basic approach [^4] The basic approach is the same as the proposed pipeline, but without video and image augmentations, regression head, and classification loss balancing. and (2) proposed strategies. We fine-tune three architectures: two transformers (MViTv2-S with different pre-training and MViTv2-B [24]) and one CNN (I3D [4] with a ResNet-50) – on four ISLR datasets. The experiment results are provided in Table 2.
@@ 5.1 Preprocessing
Input videos are resized to 300 resolution and converted to HDF5 video format. During the training stage, videos are square-padded and randomly cropped to achieve a resolution of 224 × 224. The low side of the videos is resized to 224 and square-padded during the validation and testing stages. The model input is a sequence of 32 frames sampled with a step of 2. If the video does not have enough frames for sampling, the last frame is repeated to form a complete clip. For the Slovo and SlovoExt datasets, non-mirrored signs are excluded from the horizontal flip augmentation. The WLASL signs are not being flipped, and the AUTSL signs are flipped with no exceptions because Section 6 shows the efficacy of such decisions.
@@
Table 3: Ablation study results. We employ the MViTv2-S model to assess the influence of each component by removing it from the pipeline and comparing the results. Accuracy is evaluated on the SlovoExt combination. Red values indicate differences with the entire pipeline.
@@
5.2 Training Methodology
As shown in Figure 2, the preprocessed video is initially exposed to video augmentations. One randomly selected from four modifications is applied to each batch. In contrast, “random boundary shift" is constantly affected. In experiments, “speed up" increases the video speed by 2 times, and “slow down" decreases it by 2 times. “Random drop" shortens video by 10%, while “random add" lengthen it by 30%. The random values for sign boundaries shifting are from the interval [-5, 5], except the interval [-5, 0] for the WLASL and AUTSL datasets due to their trimming by the sign boundaries.
Image augmentations modify the video batch by randomly choosing a combination of them with expertly selected probabilities and magnitudes. Note that CutMix [40] and MixUp [42] influence only transformer models in our experiments because they can degrade CNNs due to a lack of correlation between neighboring pixels [2].
@@ 5.2 Training Methodology
Training is performed until convergence on four Tesla H100s with 80GB RAM. Early stopping is triggered if the top-1 accuracy metric does not increase by at least 0.003 after 7 epochs. For the first 20 epochs, the learning rate is modified by a linear scheduler, and a cosine scheduler is used from epochs 20 to 100. Cross-entropy loss is minimized using the AdamW [28] optimizer. Other parameters are variable and can be adjusted as needed.
@@ 5.3 Results
As Table 2 shows, the proposed pipeline consistently improves top-1 accuracy across different datasets and architectures. Specifically, MViTv2-S achieves additional gains of 6.54%, 3.93%, 10.12%, and 5.76% on WLASL, AUTSL, Slovo, and SlovoExt, respectively, while I3D obtains an average improvement of 1.72% across these datasets. The ablation study (see Table 3) further underscores the effectiveness of each proposed component in the pipeline.
Moreover, this approach attains state-of-the-art performance on two benchmark ISLR datasets. On WLASL, using MViTv2-S pre-trained with MaskFeat and further pretrained on SlovoExt, the model achieves 62.89% top-1 accuracy, surpassing the NLA-SLR [47] model by 1.63%. On the Slovo dataset, this method reaches 78.21% mean accuracy, exceeding the prior best result by 14.12%. A comparable increase of 13.06% is achieved with the MViTv2-S pre-trained on K400, confirming the robustness of the proposed training strategies.
@@ 6 Ablation Study
This section estimates the impact of each part of the proposed pipeline individually. We divide all modifications into three blocks: (1) image augmentations, (2) video augmentations, and (3) additional losses. We test the necessity for each block by disconnecting it from the entire pipeline and for its parts by shutting down a particular part while not changing anything else. The ablation study uses MViTv2-S pre-trained on the MaskFeat to train on the SlovoExt combination.
@@ 6.1 Image Augmentations Necessity
The image augmentations block is additionally divided into two parts – basic image transformations and CutMix-MixUp pair – to simplify assessing their influence. Each part and combinations of them are shut down, resulting in three experiments. The first three rows in Table 3 illustrate that the metric significantly decreases by 4.06% in the absence of all image augmentations. The influence of basic image augmentation is two times less than the CutMix-MixUp one.
We also evaluate various augmentation-applying approaches, including RandAugment [7], UniformAugment [26], and random sequential applications. Table 4 indicates that RandAugment and UniformAugment approaches are not effective in the proposed pipeline.
@@ 6.2 Video Augmentations Necessity
Similar to the above subsection, we split the estimation of video augmentations necessity by disconnecting each of them separately and the whole block, providing six experiments. The most notable impact achievable with one modification is produced by the “slow down" augmentation, as without it, the top-1 accuracy drops by 0.66%.
@@ 7 Discussion
Positional Statement. Through the research of the ISLR task, we involved the All-Russian Community of Deaf experts and professional sign language interpreters. The expertise of the All-Russian Society of the Deaf was utilized at every stage of the SlovoExt dataset creation, including data collection, validation, and verification processes, as well as video tagging. We also involved deaf consultants in developing training strategies to apply Considerations to particular solutions. Additionally, some of our researchers took formal courses on RSL to enhance their knowledge of this domain.
Limitations. The regression head and IoU-balanced CE loss require sign boundary annotations that are frequently difficult to achieve. Thus, when the dataset is pre-trimmed, the “random boundary shift" augmentation can adjust the sign boundaries only inward toward the center rather than allowing shifts in both directions. We processed RGB frames and did not analyze articulation, keypoints, or depth information. Moreover, a lack of awareness regarding non-mirrored signs in sign language may lead to issues when applying flip augmentation. There are additional challenges in each sign language, and each has its specifics that must be considered when adapting the pipeline.
@@ 8 Conclusion
In this paper, we introduce a training pipeline for the ISLR models, considering the specifics of the SLR domain and the constraints of real-world usage. We demonstrate the effectiveness of applying image and video augmentations to address the issues of low data quality and varying gesturing speed. The importance of integrating temporal information into the model by the regression head combined with IoU-balanced CE loss is also presented. The developed pipeline delivers state-of-the-art results on the WLASL and Slovo datasets. Future work will extend the training strategies to CSLR and SLT tasks.
@@ Appendix B SlovoExt Creation Pipeline
We collected the self-assembled part of the SlovoExt dataset in three main stages: (1) data mining, (2) validation, and (3) annotation. Tasks were available only to those who passed the mandatory RSL exam and could thus be considered native speakers. We utilized Yandex Toloka [^7] https://platform.toloka.ai/ for the first stage and ABC Elementary [^8] https://elementary.activebc.ru for the others due to the specifics of each of the platforms and to avoid checking signers own videos. The creation pipeline of the self-assembled Russian Isolated Sign Language Dataset is briefly described below.